A sample of 114 mortgages approved during the current year showed that 37 were issued to a single-earner family or individual. The historical percentage is 29 percent. At the .05 level of significance in a right-tailed test, has the percentage of single-earner or individual mortgages risen? (a-1) H0: π ≤ .29 versus H1: π > .29. Choose the right option. Reject H0 if zcalc > 1.645 Reject H0 if zcalc < 1.645 (a-2) Calculate the test statistic. (Round intermediate calculations to 2 decimal places. Round your answer to 4 decimal places.) (a-3) The null hypothesis should be rejected (b) Is this a close decision? (c) State any assumptions that are required. nπ < 10 and n(1 − π) > 10 nπ < 10 and n(1 − π) < 10 nπ > 10 and n(1 − π) > 10 nπ > 10 and n(1 − π) < 10
Answer #1

null Hypothesis alternate Hypothesis: Ho: 0.290 Ha 0.290 for 0.05 level with right tailed test, critical value of z- Decision rule: 1.645 reject Ho if test statisticz > 1.645 sample success sample size 114.0 std error -Se p( 1-p)/n) = 0.0425 sample proportion p 0.3246 test stat- z (p-p)/Se= 0.8141

a-1) Reject H0 if zcalc > 1.645

a-2) test statistic z =0.8141 ( please try 0.82 if this comes wrong)

a-3)

nπ > 10 and n(1 − π) > 10